OpenStudy (anonymous):

A pineapple was found to have two sets of opposing spirals formed by its leaves. The clockwise spiral has x leaves. The number of leaves in the counterclockwise direction is 8 more than that in the clockwise direction. These leaf numbers are consecutive numbers in the Fibonacci sequence. Which equation is true for the number of leaves in the two spirals? A. x+8/x=1.6 B. 8/x=1.6 C. x+8/8=1.6 D. x/8=1.6

OpenStudy (anonymous):

A. x+8/x=1.6 B. 8/x=1.6 C. x+8/8=1.6 D. x/8=1.6

OpenStudy (anonymous):

I think it's A....

OpenStudy (anonymous):

1,1,2,3,5,8,13

OpenStudy (anonymous):

those are Fibonacci Numbers

OpenStudy (anonymous):

so \(x=5\) and \(x+8=13\)

OpenStudy (anonymous):

and therefore \(\frac{x+8}{x}=\frac{13}{5}=1.6\)

OpenStudy (anonymous):

i think A looks good as well

OpenStudy (anonymous):

thanks I wasn't sure :)

OpenStudy (anonymous):

https://www.youtube.com/watch?v=5NV6Rdv1a3I

OpenStudy (anonymous):

:) nice video